# Towards Imperceptible Watermarking Via Environment Illumination for Consumer Cameras

### v1

@@ 1 Introduction
To overcome this limitation, we employ temporal spectral modulation, leveraging differences in spectral sensitivity between CMOS sensors and the human visual system. Instead of modulating light intensity, we adjust the LED spectra to meet three key criteria: (1) minimal detectability by human observers, (2) maximum detectability by CMOS sensors, and (3) a white appearance equivalent to standard D65 illumination. By exploiting this spectral sensitivity disparity, we generate optical changes imperceptible to humans yet detectable by cameras, enabling low-frequency spectral modulation without visual discomfort while preserving information embedding.
[…]
Contribution To the best of our knowledge, we present the first method for optically embedding information into LED illumination sources using spectral modulation, enabling extraction from video recordings of the scene. By leveraging an optimization approach that accounts for differences in spectral sensitivity between humans and cameras, we achieve spectral variations detectable only by cameras. We validate our method’s feasibility through a user study with a prototype lighting system and various consumer cameras, demonstrating detection from videos captured across different devices and through unknown in-camera processing, while confirming that the signal remains below the human visibility threshold.
@@ 2 Related Work
This section reviews related work in visible light communication (VLC), hyperspectral lighting, and the characteristics of human and camera color perception.
Visible Light Communication VLC is a wireless communication technology that transmits data using visible light, typically emitted by LEDs [21, 1, 14, 22, 18, 15]. Unlike radio-frequency (RF) communication, VLC exploits the rapid on-off switching capabilities of LEDs, allowing them to function as both illumination and data-transmitting sources. This high-frequency flickering, imperceptible to the human eye, makes VLC promising for applications such as indoor positioning and secure data transfer in RF-sensitive environments (e.g., hospitals and airplanes).
To achieve VLC with image sensors, the low frame rate of typical cameras necessitates data transmission at relatively low frequencies near the human perceptual threshold (30–60 Hz). Several studies have explored rolling shutter image sensors, which can capture high-frequency flickering signals due to the sequential exposure of pixels across the sensor [17, 5, 12]. While this approach enables high-frequency VLC with cameras, it is impractical for one-to-many communication because variations in shutter speed and exposure time require calibration or prior knowledge of each camera’s settings. Thus, to ensure universal compatibility, it is crucial to use low-frequency, imperceptible lighting that remains undetectable to the human eye yet reliably detectable by camera sensors.
Spatial Embedding Spatial encoding techniques, such as structured light and projection mapping, embed information by projecting visible patterns onto surfaces [3, 9]. However, making these patterns imperceptible in everyday environments is challenging. Even if achieved, these methods require precise calibration of the projection surface, making large-scale deployment costly and impractical. In contrast, our approach eliminates the need for spatial alignment by embedding information directly into ambient lighting through temporal spectral modulation, ensuring seamless integration into real-world settings.
@@
Hyperspectral Illumination Recent advances in narrowband LED technology have enabled the construction of hyperspectral lighting systems by combining multiple LEDs [24]. As shown in Fig. 2, the spectrum of a narrowband LED lighting system can be customized by adjusting the emission intensities of individual LEDs, which combine as a weighted sum of each LED’s spectral profile. Traditionally, controllable hyperspectral lighting required specialized light sources, typically involving diffraction gratings or spatial light modulators [25]. In contrast, the multi-LED approach offers a simpler and more accessible way to design lighting with tailored spectral characteristics, enabling customized spectral profiles.
Human Color Perception and Metamerism Human color perception is governed by the trichromatic theory, which states that color vision relies on three types of photoreceptor cells in the retina, each sensitive to different regions of the visible spectrum. These sensitivities, often represented by the CIE XYZ color matching functions[6, 23], serve as the foundation for standard color spaces that model human vision [2, 11]. However, metamerism allows different spectral power distributions (SPDs) to produce identical color sensations for human observers [7, 20]. This occurs because the brain interprets color based on the combined responses of the three photoreceptor types, meaning distinct light spectra can elicit the same perceptual response.
Figure 3: Proposed optimization pipeline. The objective functions, shown from top to bottom, ensure human imperceptibility, camera detectability, and natural white light for everyday use. These functions guide the optimization of LED intensity variables to achieve the desired lighting characteristics.
Comparing Camera and Human Spectral Sensitivity Consumer cameras approximate human color perception using three primary color filters. However, achieving an exact match with human visual sensitivity is challenging due to technical and design constraints. Factors such as filter material properties, manufacturing costs, and brand-specific design choices create variations in camera spectral sensitivity. To improve visual accuracy, manufacturers apply post-processing techniques to adjust colors, making images appear more natural to human viewers [10, 8]. Linear transformations are commonly used to align the camera’s color space with the human XYZ space, but these adjustments are inherently limited. As a result, camera sensors still differ from human vision in their spectral response [13].
@@ 3.1 Spectral Optimization
Human Perceptibility
The perceptibility of color differences to humans has been extensively studied in color difference research and is quantified using DeltaE2000[4, 19]. For our method, we use DeltaE2000($\Delta_{\mathrm{E}}$) to ensure that the reflected light from all objects remains imperceptible to humans. To meet this criterion, we define the following evaluation function:
@@
A difference of $\Delta_{\mathrm{E}}\leq 2$ is generally imperceptible to untrained observers when colors are viewed side by side [4, 19]. However, in practical applications, the human eye is more sensitive to low-frequency flicker, making it more likely to detect subtle differences when colors alternate over time, even if these differences are imperceptible in a static, side-by-side comparison. Therefore, the evaluation function is minimized under this constraint, provided that the requirements of the other loss functions are satisfied.
Camera Detectability To ensure the watermark remains imperceptible, it is crucial that the flicker does not produce excessive pixel differences that could be easily noticeable to viewers. If the flicker in the captured video is too prominent, it may significantly degrade visual quality, making the video unsuitable for typical viewing. The following evaluation function is designed to ensure adequate detectability across all spectral channels while controlling pixel differences for multiple camera models:
@@
Natural Illumination Another criterion for this method is that the light must maintain a neutral white appearance with minimal spectral skew, ensuring suitability for everyday use. Color rendering is conventionally evaluated using the Color Rendering Index (CRI), which measures color fidelity based on standardized samples. CRI values range from 0 to 100, with higher values indicating more natural color rendering. Common household LEDs and bulbs typically have CRI values around 60, which is generally sufficient for everyday applications.
Our approach targets the standard D65 illumination (i.e., spectra that appears as daylight) and aim for a CRI value exceeding 60, calculated according to the CRI standard, which uses eight specific color samples to evaluate general lighting color rendering. The CRI value is derived using the formula $\text{CRI}=100-4.6\times\Delta_{\mathrm{E}}$ , where $\Delta_{\mathrm{E}}$ represents the average color difference between each sample illuminated by the test light and the reference (D65). Optimization is performed using the following evaluation function:
$\mathcal{L}_{\mathrm{w}}$ $=\sum_{i}\mathrm{ReLU}\left(\Delta_{1}-\tau_{\mathrm{w}}\right)+\mathrm{ReLU}\left(\Delta_{2}-\tau_{\mathrm{w}}\right)$ (6)
$\Delta_{1}$ $=\Delta_{\mathrm{E}}\left(\boldsymbol{s}_{\mathrm{h}}\left(\boldsymbol{R}_{\mathrm{o}}[i]\odot\boldsymbol{L}_{\mathrm{ref}}\right),\boldsymbol{s}_{\mathrm{h}}\left(\boldsymbol{R}_{\mathrm{o}}[i]\odot\boldsymbol{L}_{1}\right)\right)$ (7)
$\Delta_{2}$ $=\Delta_{\mathrm{E}}\left(\boldsymbol{s}_{\mathrm{h}}\left(\boldsymbol{R}_{\mathrm{o}}[i]\odot\boldsymbol{L}_{\mathrm{ref}}\right),\boldsymbol{s}_{\mathrm{h}}\left(\boldsymbol{R}_{\mathrm{o}}[i]\odot\boldsymbol{L}_{2}\right)\right),$ (8)
where $\boldsymbol{L}_{\mathrm{ref}}$ represents the D65 reference spectrum and $\tau_{\mathrm{w}}$ is a threshold that ensures achieving a CRI value of 60 or higher. We set this threshold to 40/4.6 , equivalent to a CRI value of 60.
Weighted Optimization The optimal two spectra are designed by adjusting and minimizing these three losses as follows:
@@
Reparameterization for Hyperspectral Illumination As shown in Fig 2, hyperspectral illumination can be reproduced as a weighted sum of multiple LEDs. Therefore, rather than directly optimizing $\boldsymbol{L}_{1}$ and $\boldsymbol{L}_{2}$, we optimize the intensity of each LED. Let each LED’s spectral profile be denoted as $\boldsymbol{\Lambda}i$ and their respective intensities be $x_{i}$ and $y_{i}$. The spectral compositions of the lighting follow:
$\boldsymbol{L}_{1}=\sum_{i=1}^{N}x_{i}\boldsymbol{\Lambda}_{i},\quad\boldsymbol{L}_{2}=\sum_{i=1}^{N}y_{i}\boldsymbol{\Lambda}_{i},$ (10)
where $N$ is the number of LEDs. In this manner, reparameterization shifts the optimization target to $\boldsymbol{x}=[x_{0},\cdots,x_{n}]$ and $\boldsymbol{y}=[y_{0},\cdots,y_{n}]$.
@@ 3.2 Detection of Flicker
The following briefly explains the detection algorithm used by our approach. The supplementary materials provide a more detailed explanation.
Frames per Second (fps) and Bits per Second (bps) In our setup, lighting fluctuations are encoded as binary values, with frame-to-frame pixel intensity changes representing bits (1 for a change, 0 for no change). This differential encoding generates a bit sequence based on frame differences across the image. According to Shannon’s theorem, a recording rate of $n$ fps ideally enables communication at $n$ bps. However, real cameras have finite exposure times, which can introduce inaccuracies if phase shifts occur between illumination changes and the camera’s sampling phase. To mitigate this, we set the recording rate to $2n$ fps, ensuring synchronization and achieving $n$ bps.
Signal in Dark Regions In differential encoding, frame-to-frame differences may be lost in dark regions due to 8-bit quantization, as pixel fluctuations are often too small to detect accurately. To improve detectability, we set the camera detectability loss so that the average pixel fluctuation is approximately 1. Pixels with values consistently below 128 across all channels and time are masked to prevent quantization loss.
Normalization To ensure detection accuracy despite differences in scene brightness due to camera settings like exposure time and white balance, we normalize pixel intensity values across all channels, setting the mean intensity to 1. This normalization standardizes variations, enabling consistent results across different camera configurations.
Temporal Differential Calculation With normalized data, we compute differences along the time axis to capture frame-to-frame changes representing encoded binary values. This is achieved by subtracting each frame from the previous one. At this stage, differential information is obtained at a doubled sampling rate to mitigate phase shift effects. For final processing, we return to the signal’s original sampling rate by downsampling, selecting the phase that best captures the lighting fluctuations to address potential misalignment.
Signal Binarization Finally, we compute the average brightness changes across the image for each time frame to capture fluctuations over time. To extract a binary signal that adapts to the scene’s characteristics, we use an adaptive thresholding approach. We begin with a static threshold to estimate the average brightness fluctuation level when significant changes occur. This serves as a baseline for assessing typical fluctuation intensity. Based on this baseline, we then define a dynamic threshold set to half of the average fluctuation, ensuring consistent identification of significant brightness changes as binary signals.
@@ 4.1 Spectral Optimization
For the prototype lighting setup, we use ten types of LEDs with different wavelengths, as shown in Fig. 4. LED brightness is modulated by adjusting the current through a constant-current chip using pulse-width modulation (PWM) via a Raspberry Pi. During optimization, the duty cycle of each LED serves as the adjustable variable for brightness control.
Camera sensor characteristics for the optimization are taken from a publicly available dataset of spectral sensitivities for 28 cameras [13]. The optimization problem is solved using the Adam optimizer [16] with a learning rate of $0.01$, over 5000. The hyperparameters for optimization, denoted as $\boldsymbol{w}_{h},\boldsymbol{w}_{c},\boldsymbol{w}_{w}$, are determined through grid search. The hyperparameters used in this case are $\boldsymbol{w}_{h}=0.15,\boldsymbol{w}_{c}=0.05,\boldsymbol{w}_{w}=0.8$.
The final spectra obtained through this optimization process are shown in Fig. 5. The resulting average $\Delta_{\mathrm{E}}$ of $0.2839$ indicates minimal color difference, remaining indistinguishable to the human eye.

### v2

@@ 1 Introduction
To overcome this limitation, we employ temporal spectral modulation, leveraging differences in spectral sensitivity between CMOS sensors and the human visual system [jiang2013space]. Instead of modulating light intensity, we adjust the LED spectra to meet three key criteria: (1) minimal detectability by human observers, (2) maximum detectability by CMOS sensors, and (3) a white appearance equivalent to standard D65 illumination. By exploiting this spectral sensitivity disparity, we generate optical changes imperceptible to humans yet detectable by cameras, enabling low-frequency spectral modulation without visual discomfort while preserving information embedding.
Although the system transmits data at a relatively low rate of 15 bits per second, it can embed up to 128 bits within a 10-second video clip. This capacity is sufficient to encode essential metadata, such as location information, recording permissions, and timestamps–key details that support privacy protection and content verification.
Contributions To the best of our knowledge, we present the first method for optically embedding information into LED illumination sources using spectral modulation, enabling extraction from video recordings of the scene. By leveraging an optimization approach that accounts for differences in spectral sensitivity between humans and cameras, we achieve spectral variations detectable only by cameras. We validate our method’s feasibility through a user study with a prototype lighting system and various consumer cameras, demonstrating detection from videos captured across different devices and through unknown in-camera processing, while confirming that the signal remains below the human visibility threshold. Beyond watermarking, our approach may serve as a human-invisible, camera-centric active cue for broader vision tasks—for example, near-infrared “dark-flash” illumination that aids geometry/appearance estimation [darkflash_iccv21]—with related explorations indicating similar potential for depth sensing and for hyperspectral (and HS+3D) reconstruction using commodity RGB cameras [activestereonet_eccv18, park_iccv07, shin_cvpr24, shin_cvpr25].
@@ 2 Related Work
This section reviews related work in visible light communication (VLC), spatial watermarking, hyperspectral lighting, and the characteristics of human and camera color perception.
Visible Light Communication (VLC). VLC is a wireless communication technology that transmits data using visible light, typically emitted by LEDs [tanaka2000wireless, ieee2011ieee, karunatilaka2015led, tsonev2014light, sevincer2013lightnets, khalighi2014survey]. Unlike radio-frequency (RF) communication, VLC exploits the rapid on-off switching capabilities of LEDs, allowing them to function as both illumination and data-transmitting sources. This high-frequency flickering, imperceptible to the human eye, makes VLC promising for applications such as indoor positioning and secure data transfer in RF-sensitive environments (e.g., hospitals and airplanes).
To achieve VLC with image sensors, the low frame rate of typical cameras necessitates data transmission at relatively low frequencies near the human perceptual threshold (30–60 Hz). Several studies have explored rolling shutter image sensors, which can capture high-frequency flickering signals due to the sequential exposure of pixels across the sensor [Liang, Chow2015, Ji]. While this approach enables high-frequency VLC with cameras, it is impractical for one-to-many communication because variations in shutter speed and exposure time require calibration or prior knowledge of each camera’s settings. Thus, to ensure universal compatibility, it is crucial to use low-frequency, imperceptible lighting that remains undetectable to the human eye yet reliably detectable by camera sensors.
Spatial Embedding. Recent CVPR/ICCV work has rapidly advanced invisible digital watermarking for validating generative content [zhang_cvpr25_omniguard, bui_iccv25_trustmark]. Similarly, in the spatial domain, information can be embedded by projecting patterns using structured light or projection mapping [Amiri, Geng_11]. Although these spatial methods can embed more information through spatial addressing and multiplexing, they usually require pre-mapping the scene and precise alignment between the projector and surfaces, which makes them difficult to deploy and maintain in large or changing environments. In addition, making projected cues truly imperceptible in everyday settings remains challenging.
@@
Hyperspectral Illumination. Recent advances in narrowband LED technology have enabled the construction of hyperspectral lighting systems by combining multiple LEDs [zhao2020narrow]. As shown in Fig. 2, the spectrum of a narrowband LED lighting system can be customized by adjusting the emission intensities of individual LEDs, which combine as a weighted sum of each LED’s spectral profile. Traditionally, controllable hyperspectral lighting required specialized light sources, typically involving diffraction gratings or spatial light modulators [zheng2015separating]. In contrast, the multi-LED approach offers a simpler and more accessible way to design lighting with tailored spectral characteristics, enabling customized spectral profiles.
Human Color Perception and Metamerism. Human color perception is governed by the trichromatic theory, which states that color vision relies on three types of photoreceptor cells in the retina, each sensitive to different regions of the visible spectrum. These sensitivities, often represented by the CIE XYZ color matching functions [cie1931proceedings, ws82], serve as the foundation for standard color spaces that model human vision [agoston2013color, H84]. However, metamerism allows different spectral power distributions (SPDs) to produce identical color sensations for human observers [fairchild2013color, SSBR07]. This occurs because the brain interprets color based on the combined responses of the three photoreceptor types, meaning distinct light spectra can elicit the same perceptual response.
Figure 3: Proposed optimization pipeline. The objective functions ensure human imperceptibility, camera detectability, and natural white light for everyday use. These functions guide the optimization of LED intensity variables to achieve the desired lighting characteristics. Dark blue and dark green arrows indicate the forward pass for the two illuminations, while red arrows denote the backpropagation path used to optimize the LED intensities.
Comparing Camera and Human Spectral Sensitivity. Consumer cameras approximate human color perception using three primary color filters. However, achieving an exact match with human visual sensitivity is challenging due to technical and design constraints. Factors such as filter material properties, manufacturing costs, and brand-specific design choices create variations in camera spectral sensitivity. Due to differences in spectral sensitivities, cameras apply color space mappings to convert sensor-specific RGB values to a human-perceptual color space  [HLR01, hakki_color, kim2025ccmnet]. Linear transformations are commonly used to align the camera’s color space with the human XYZ space, but these adjustments are inherently limited. As a result, camera sensors still differ from human vision in their spectral response [jiang2013space].
@@ 3.1 Spectral Optimization
Human Perceptibility. The perceptibility of color differences to humans has been extensively studied in color difference research and is quantified using DeltaE2000 [brainard2003color, stokes1992]. For our method, we use DeltaE2000($\Delta_{\mathrm{E}}$) to ensure that the reflected light from all objects remains imperceptible to humans. To meet this criterion, we define the following evaluation function:
@@
A difference of $\Delta_{\mathrm{E}}\leq 2$ is generally imperceptible to untrained observers when colors are viewed side by side [brainard2003color, stokes1992]. However, in practical applications, the human eye is more sensitive to low-frequency flicker, making it more likely to detect subtle differences when colors alternate over time, even if these differences are imperceptible in a static, side-by-side comparison. Therefore, the evaluation function is minimized under this constraint, provided that the requirements of the other loss functions are satisfied.
Camera Detectability. To ensure the watermark remains imperceptible, it is crucial that the flicker does not produce excessive pixel differences that could be easily noticeable to viewers. If the flicker in the captured video is too prominent, it may significantly degrade visual quality, making the video unsuitable for typical viewing. The following evaluation function is designed to ensure adequate detectability across all spectral channels while controlling pixel differences for multiple camera models:
@@
Natural Illumination. Another criterion for this method is that the light must maintain a neutral white appearance with minimal spectral skew, ensuring suitability for everyday use. Color rendering is conventionally evaluated using the Color Rendering Index (CRI), which measures color fidelity based on standardized samples. CRI values range from 0 to 100, with higher values indicating more natural color rendering. Common household LEDs and bulbs typically have CRI values around 60, which is generally sufficient for everyday applications.
Our approach targets the standard D65 illumination (i.e., spectra that appear as daylight) and enforces a CRI of at least 60 per the CRI standard (eight test samples). We define $\mathrm{CRI}=100-4.6\,\Delta E$, where $\Delta E$ is the average color difference to D65. To ensure $\mathrm{CRI}\geq 60$, we minimize a hinge penalty on $\Delta E$ (Eq. 6).
$\mathcal{L}_{\mathrm{w}}$ $=\sum_{i=1}^{8}\operatorname{ReLU}\!\left(\Delta_{1,i}-\tau_{\mathrm{w}}\right)+\operatorname{ReLU}\!\left(\Delta_{2,i}-\tau_{\mathrm{w}}\right)$ (6)
$\Delta_{1,i}$ $=\Delta E\!\big(\boldsymbol{s}_{\mathrm{h}}(\boldsymbol{R}_{\mathrm{o}}[i]\odot\boldsymbol{L}_{\mathrm{ref}}),~\boldsymbol{s}_{\mathrm{h}}(\boldsymbol{R}_{\mathrm{o}}[i]\odot\boldsymbol{L}_{1})\big)$ (7)
$\Delta_{2,i}$ $=\Delta E\!\big(\boldsymbol{s}_{\mathrm{h}}(\boldsymbol{R}_{\mathrm{o}}[i]\odot\boldsymbol{L}_{\mathrm{ref}}),~\boldsymbol{s}_{\mathrm{h}}(\boldsymbol{R}_{\mathrm{o}}[i]\odot\boldsymbol{L}_{2})\big),$ (8)
where $\boldsymbol{L}_{\mathrm{ref}}$ is the D65 reference spectrum and $\tau_{\mathrm{w}}=40/4.6$ enforces $\mathrm{CRI}\geq 60$.
Weighted Optimization. The optimal two spectra are designed by adjusting and minimizing these three losses as follows:
@@
Reparameterization for Hyperspectral Illumination. As shown in Fig. 2, hyperspectral illumination can be reproduced as a weighted sum of multiple LEDs. Therefore, rather than directly optimizing $\boldsymbol{L}_{1}$ and $\boldsymbol{L}_{2}$, we optimize the intensity of each LED. Let each LED’s spectral profile be denoted as $\boldsymbol{\Lambda}i$ and their respective intensities be $x_{i}$ and $y_{i}$. The spectral compositions of the lighting follow:
$\boldsymbol{L}_{1}=\sum_{i=1}^{m}x_{i}\boldsymbol{\Lambda}_{i},\quad\boldsymbol{L}_{2}=\sum_{i=1}^{m}y_{i}\boldsymbol{\Lambda}_{i},$ (10)
where $m$ is the number of LEDs. In this manner, reparameterization shifts the optimization target to $\boldsymbol{x}=[x_{0},\cdots,x_{m}]$ and $\boldsymbol{y}=[y_{0},\cdots,y_{m}]$.
@@ 3.2 Detection of Flicker
We now outline the detection algorithm. The supplementary material provides a more detailed explanation.
Frames per Second (fps) and Bits per Second (bps). In our setup, lighting fluctuations are encoded as binary values, with frame-to-frame pixel intensity changes representing bits (1 for a change, 0 for no change). This differential encoding generates a bit sequence based on frame differences across the image. According to Shannon’s theorem, a recording rate of $n$ fps ideally enables communication at $n$ bps. However, real cameras have finite exposure times, which can introduce inaccuracies if phase shifts occur between illumination changes and the camera’s sampling phase. To mitigate this, we set the recording rate to $2n$ fps, ensuring synchronization and achieving $n$ bps.
Signal in Dark Regions. In dark regions, 8-bit quantization can erase inter-frame differences, as pixel fluctuations are often too small to detect accurately. To improve detectability, we set the camera detectability loss so that the average pixel fluctuation is approximately 1. Pixels with values consistently below 128 across all channels and time are masked to prevent quantization loss.
Normalization. To ensure detection accuracy despite differences in scene brightness due to camera settings like exposure time and white balance, we normalize pixel intensity values across all channels, setting the mean intensity to 1. This normalization standardizes variations, enabling consistent results across different camera configurations.
Temporal Differential Calculation. With normalized data, we compute differences along the time axis to capture frame-to-frame changes representing encoded binary values. This is achieved by subtracting each frame from the previous one. At this stage, differential information is obtained at $2\timessampling$ rate to mitigate phase shifts. For final processing, we return to the signal’s original sampling rate by downsampling, selecting the phase that best captures the lighting fluctuations to address potential misalignment.
Signal Binarization. Finally, we compute the average brightness changes across the image for each time frame to capture fluctuations over time. To extract a binary signal that adapts to the scene’s characteristics, we use an adaptive thresholding approach. We begin with a static threshold to estimate the average brightness fluctuation level when significant changes occur. This serves as a baseline for assessing typical fluctuation intensity. Based on this baseline, we then define a dynamic threshold set to half of the average fluctuation, ensuring consistent identification of significant brightness changes as binary signals.
@@ 4.1 Spectral Optimization
For the prototype lighting setup, we use ten types of LEDs with different wavelengths, as shown in Fig. 4. LED brightness is modulated by adjusting the current through a constant-current chip using pulse-width modulation via a Raspberry Pi. During optimization, the duty cycle of each LED serves as the adjustable variable for brightness control.
Camera sensor characteristics for the optimization are taken from a publicly available dataset of spectral sensitivities for 28 cameras [jiang2013space]. The optimization problem is solved using the Adam optimizer [kingma2014adam] with a learning rate of $0.01$, over 5000 iterations. The hyperparameters for optimization, denoted as $\boldsymbol{w}_{h},\boldsymbol{w}_{c},\boldsymbol{w}_{w}$, are determined through grid search. The hyperparameters used in this case are $\boldsymbol{w}_{h}=0.15,\boldsymbol{w}_{c}=0.05,\boldsymbol{w}_{w}=0.8$.
The final spectra obtained through this optimization process are shown in Fig. 5. The optimized pair yields a mean $\Delta E=0.2839$, i.e., imperceptible.